The system shown can undergo vertical motion only, and both springs are unstretched
when WD0. Determine the deflection ?in terms of W, k1,andk2

The system shown can undergo vertical motion only,
Answer #1

At the equilibrium, the net force on the mass is null.

F_{\mathrm{net}}=0

then,

F_{1}+F_{2}-W=0

where,

  • F_{1} is the magnitude of force exerted by the spring 1

F_{1}=K_{1}\delta

  • F_{2} is the magnitude of the force exerted by the spring 2

F_{2}=K_{2}\delta

then,

F_{1}+F_{2}-W=0

K_{1}\delta+K_{2}\delta-W=0

according to the above expression, the deflection must be

\left( K_{1}+K_{2} \right)\delta=W

\delta=\frac{W}{K_{1}+K_{2}}

so, when

  • W=0.

\delta=\frac{0}{K_{1}+K_{2}}=0

if the weight is null there is not deflection

  • K_{1}=0

\delta=\frac{W}{0+K_{2}}=\frac{W}{K_{2}}

  • K_{2}=0

\delta=\frac{W}{K_{1}+0}=\frac{W}{K_{1}}

  • K_{1}\rightarrow \infty

K_{1}+K_{2}\rightarrow \infty

then

\delta=\frac{W}{K_{1}+K_{2}} \rightarrow 0

  • K_{2}\rightarrow \infty

K_{1}+K_{2}\rightarrow \infty

then

\delta=\frac{W}{K_{1}+K_{2}} \rightarrow 0